Okay, sorry, I have another question. These radicals are killing me!
sqrt ( x-3 ) + sqrt ( x+2 ) = 5
Really hard for a student do when there's not ONE example to turn to look for help in the book. Sheesh. Thank you again in advance!

Answer by scott8148(6628)

squaring ___ (x - 3) + 2 sqrt[(x - 3)(x + 2)] + (x + 2) = 25

2 sqrt(x^2 - x - 6) = 26 - 2x ___ sqrt(x^2 - x - 6) = 13 - x

squaring (again) ___ x^2 - x - 6 = 169 - 26x + x^2